OpenStudy (anonymous):

Can you guys help me by checking my work? Simplify: 10radical2x^3 - xradical98x + radical50x^3. My answer: 8radical2x. Is that correct?

OpenStudy (lgbasallote):

\[10\sqrt{2x^3} - x\sqrt{98x} + \sqrt{50x^3}\] is that the question?

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

\[\sqrt{2x^3} = \sqrt{2 \times x \times x^2} = x\sqrt{2x}\] \[\sqrt{98x} = \sqrt{2 \times 49 \times x} = 7\sqrt{2x}\] \[\sqrt{50x^3} = \sqrt{2 \times 25 \times x^2 \times x} = 5x \sqrt{2x}\] so if we simplify we have \[\large 10x \sqrt{2x} - 7x \sqrt{2x} + 5x\sqrt{2x}\] \[\large \implies 8x\sqrt{2x}\] \[\huge \checkmark\] you're right

OpenStudy (anonymous):

Thanks :)